# Imaging pulsar census of the Galactic Plane using MWA VCS data

### v1

@@ 4.5 Spectral energy distributions
With only two flux density measurements, two-point spectral index can be calculated analytically (without fitting). Therefore, in order to increase the reliability of the spectral fits produced, the software, pulsar_spectra, requires at least three data points for a simple power law fit and four or more data points for other spectral models. The software does not produce any fits if this requirement is not satisfied. Taking this threshold into consideration, out of the 66 pulsars, 63 had sufficient number of flux density measurements in the literature to produce the spectral fits without the addition of our work, one did not have enough measurements to fit the model without our data points and two did not have a sufficient number of data points to create a spectral fit even after the addition of our work. The lack of measurement data at frequencies below 400 MHz shows the need for more low-frequency campaigns to fill the gap in the literature. The reliable flux density measurement of pulsars with the MWA below 300 MHz will be useful in better spectral modelling of the pulsars. This is exemplified by the fact that 15 out of the 66 pulsars that were detected in Stokes I images changed the spectral models after the addition of our measurements as shown in Table 5. The table also shows the Akaike Information Criteria (AIC), which is a measure of how much information the model retains without over-fitting. The model that results in the lowest AIC is the most likely to be a sufficiently accurate model, which does not over-fit the data. The AIC of the model before and after the addition of the flux density from this work is given in Table 5. We note that the AIC value may not be ideal to compare the accuracy of the fit before and after the addition of our points. This is because AIC is mainly aimed at finding the ”minimal model” which describes the data without over-fitting sufficiently well. This demonstrates lack of low-frequency flux densities in the literature causes large uncertainties and volatility of the spectral modelling (i.e. addition of a single data point can change the fitted spectral model).
PSRJ Name | Original model | New model | AIC(old) | AIC(new) |
J0820-1350 | LFTO | DTOS | 231.4 | 235.5 |
J0820-4114 | BPL | LFTO | 89.9 | 89.2 |
J0907-5157 | LFTO | BPL | 43.2 | 42.9 |
J1121-5444 | HFCO | LFTO | 25.1 | 24.8 |
J1312-5402 | SPL | LFTO | 24.0 | 21.8 |
J1456-6843 | LFTO | BPL | 88.6 | 100.0 |
J1709-1640 | HFCO | BPL | 157.9 | 156.6 |
J1752-2806 | DTOS | LFTO | 740.4 | 893.4 |
J1913-0440 | LFTO | BPL | 226.3 | 206.4 |
J1932+1059 | LFTO | DTOS | 166.2 | 169.9 |
J2022+2854 | LFTO | BPL | 141.4 | 137.6 |
J2048-1616 | LFTO | HFCO | 323.6 | 323.8 |
J1600-5044 | LFTO | BPL | 45.7 | 42.4 |
J1807-2715 | SPL | HFCO | 24.6 | 20.4 |
J1827-0958 | SPL | HFCO | 20.3 | 13.0 |

### v2

@@ 4.5 Spectral energy distributions
With only two flux density measurements, two-point spectral index can be calculated analytically (without fitting). Therefore, in order to increase the reliability of the spectral fits produced, the software, pulsar_spectra, requires at least three data points for a simple power law fit and four or more data points for other spectral models. The software does not produce any fits if this requirement is not satisfied. Taking this threshold into consideration, out of the 66 pulsars, 63 had sufficient number of flux density measurements in the literature to produce the spectral fits without the addition of our work, one did not have enough measurements to fit the model without our data points and two did not have a sufficient number of data points to create a spectral fit even after the addition of our work. The lack of measurement data at frequencies below 400 MHz shows the need for more low-frequency campaigns to fill the gap in the literature. The reliable flux density measurement of pulsars with the MWA below 300 MHz will be useful in better spectral modelling of the pulsars. This is exemplified by the fact that 15 out of the 66 pulsars that were detected in Stokes I images changed the spectral models after the addition of our measurements as shown in Table 5. The table also shows the Akaike Information Criteria (AIC), which is a measure of how much information the model retains without over-fitting. The model that results in the lowest AIC is the most likely to be a sufficiently accurate model, which does not over-fit the data. Three different values of AIC is given: one which corresponds to model which fits the literature data (), one for the model that fits the data after the addition of data from this work () and one for the previous model when fit to the new data set (). However, it is important to note that AIC comparisons should only be done between different models applied to the same datasets. This means that it makes sense to compare with , but it does not make sense to compare and . The lowest AIC when comparing and gives us the model that best fits the data after the addition of the flux densities from this work. For ten of the pulsars the new model is the preferred one, whereas the old model is preferred for five of the pulsars. The change in the value of AIC ranges from as small as 1 to as large as 85. Any reduction in AIC is a meaningful improvement in the model and therefore even a small change is still significant. A rule of thumb is that a differences in AIC of the order of 20% suggest preference to the model with a lower AIC value, whereas smaller differences can be treated as indication but are not really able to distinguish different models (ref:aic). Table 5 shows that there are some cases like for PSRs J1312-5402 or J1807-2715 where the new model is much more preferred than the old model i.e is much lower then . There is also one case for PSR J1456-6843 where the old model is much more preferred. However, in many cases the differences between and are smaller than 20% and should be treated as an indication only. Further improvement can be achieved by incorporating more data covering sparsely populated frequency ranges. This demonstrates lack of low-frequency flux densities in the literature causes large uncertainties and volatility of the spectral modelling (i.e. addition of a single data point can change the fitted spectral model).
| PSRJ Name | Old model | 333old model old data | New model | 444new model new data | 555old model new data |
| J0820-1350 | LFTO | 231.4 | DTOS | 249.2 | 247.6 |
| J0820-4114 | BPL | 89.9 | LFTO | 110.6 | 115 |
| J0907-5157 | LFTO | 43.2 | BPL | 42.2 | 44 |
| J1121-5444 | HFCO | 25.1 | LFTO | 21.8 | 25.4 |
| J1312-5402 | SPL | 24.0 | LFTO | 21.8 | 107.6 |
| J1456-6843 | LFTO | 88.6 | BPL | 122.8 | 96.7 |
| J1709-1640 | HFCO | 157.9 | BPL | 107.7 | 118.4 |
| J1752-2806 | DTOS | 740.4 | LFTO | 143.5 | 131.5 |
| J1913-0440 | LFTO | 226.3 | BPL | 228.8 | 282.8 |
| J1932+1059 | LFTO | 166.2 | DTOS | 187 | 158.2 |
| J2022+2854 | LFTO | 141.4 | BPL | 105.2 | 106.2 |
| J2048-1616 | LFTO | 323.6 | HFCO | 295.2 | 185.1 |
| J1600-5044 | LFTO | 45.7 | BPL | 41.8 | 46.8 |
| J1807-2715 | SPL | 24.6 | HFCO | 24.3 | 45.4 |
| J1827-0958 | SPL | 20.3 | HFCO | 20 | 29.1 |